# Moduli stacks of Higgs bundles on stable curves

### v2

@@ 5 Completeness of the Hitchin map
Notice now:
  1. 1.
    if the spectrum $T$ of the discrete valuation ring is not faithfully flat over $S$ then the surface $\mathcal{X}\times_{S}T$ is isomorphic to the product $X_{0}\times T$
  2. 2.
    if the spectrum $T$ of the discrete valuation ring is faithfully flat over $S$ then the surface $\mathcal{X}\times_{S}T$ is isomorphic to a normal surface with an isolated singularity of type $\frac{\mathbb{k}[[x,y,t]]}{xy-t^{d}}$.
In the second case, it is a routine check that [15, Proposition 6] holds for the normal surface. Indeed, any vector bundle $\mathcal{E}^{o}$ on the generic fibre $\mathcal{X}\times_{S}T^{o}=\mathcal{X}\times_{S}\operatorname{Spec}(K)$ can be extended to the surface $\mathcal{X}\times_{S}T$ as a torsion-free sheaf $\mathcal{F}$. Write $i:\mathcal{X}\times_{S}T^{o}\hookrightarrow\mathcal{X}\times_{S}T$ for the natural inclusion, and suppose we are given a Higgs field $\phi^{o}:\mathcal{E}^{o}\to\mathcal{E}^{o}\otimes i^{*}\omega_{\mathcal{X}_{T}/T}$ on $\mathcal{E}^{o}$. To extend $\phi^{o}$ over all of $\mathcal{X}\times_{S}T$ consider the following diagram
${\mathcal{F}}$${i_{*}\mathcal{E}^{o}}$${i_{*}\mathcal{E}^{o}/{\mathcal{F}}}$00{0}00{0}${\mathcal{F}\otimes\omega_{\mathcal{X}_{T}/T}}$${i_{*}(\mathcal{E}^{o}\otimes i^{*}\omega_{\mathcal{X}_{T}/T})}$${(i_{*}\mathcal{E}^{o}/{\mathcal{F}})\otimes\omega_{\mathcal{X}_{T}/T}}$00{0}$\scriptstyle{i_{*}\phi^{o}}$ (5.6)
We wish to prove the existence of a Higgs field $\mathcal{F}\to\mathcal{F}\otimes\omega_{\mathcal{X}_{T}/T}$ making the diagram commute. Using [24, Lemma 6.5], we can extend the Higgs field on $\mathcal{F}$ except at the node. Therefore, the composite map $\mathcal{F}\longrightarrow(i_{*}\mathcal{E}^{o}/\mathcal{F})\otimes\omega_{\mathcal{X}_{T}/T}$ vanishes at the generic point of the closed fibre $X_{0}$. Therefore, the obstruction to the existence of the Higgs field we seek lives in the torsion subsheaf of $i_{*}\mathcal{E}^{o}/\mathcal{F}$, when $i_{*}\mathcal{E}^{o}/\mathcal{F}$ is viewed as a sheaf on its support $X_{0}$. Therefore, if we show that the sheaf $i_{*}\mathcal{E}^{o}/\mathcal{F}$ is pure, i.e. is torsion-free on its support $X_{0}$, then we will show that the induced Higgs field on $\mathcal{F}$ extends everywhere. To see this we restrict the upper exact sequence to the closed fibre $X_{0}$. We get the exact sequence
$0\longrightarrow\mathcal{T}or^{1}_{{}_{\mathcal{X}_{T}}}(i_{*}\mathcal{E}^{o}/{\mathcal{F}},\mathcal{O}_{X_{0}})\longrightarrow\mathcal{F}|_{X_{0}}\longrightarrow(i_{*}\mathcal{E}^{o})|_{X_{0}}\longrightarrow i_{*}\mathcal{E}^{o}/{\mathcal{F}}\longrightarrow 0$ (5.7)
However, $\mathcal{T}or^{1}_{{\mathcal{X}_{T}}}(i_{*}\mathcal{E}^{o}/{\mathcal{F}},\mathcal{O}_{X_{0}})\cong(i_{*}\mathcal{E}^{o}/{\mathcal{F}})\otimes\mathcal{O}_{{}_{\mathcal{X}_{T}}}(-X_{0})$. Since $\mathcal{O}_{{}_{\mathcal{X}_{T}}}(-X_{0})$ is a locally-free sheaf, $i_{*}\mathcal{E}^{o}/\mathcal{F}$ is pure if and only if $\mathcal{T}or^{1}_{{\mathcal{X}_{T}}}(i_{*}\mathcal{E}^{o}/{\mathcal{F}},\mathcal{O}_{X_{0}})$ is pure. But $\mathcal{T}or^{1}_{{\mathcal{X}_{T}}}(i_{*}\mathcal{E}^{o}/{\mathcal{F}},\mathcal{O}_{X_{0}})$, viewed as a sheaf on its support, is a subsheaf of $\mathcal{F}|_{X_{0}}$, therefore as a sheaf on $X_{0}$ it is torsion-free. This completes the argument in the second case.
In the first case, the surface $\mathcal{X}_{T}\cong X_{0}\times T$. Therefore, the normalisation of $\mathcal{X}_{T}$ is isomorphic to $\widetilde{X_{0}}\times T$. Let us denote the normalisation morphism by $f:\widetilde{X_{0}}\to X_{0}$. From [10, Proposition 3.2], it follows that we can extend the pullback of the generic Higgs bundle $((f\times\mathrm{id}_{T^{o}})^{*}\mathcal{E}^{o},(f\times\mathrm{id}_{T^{o}})^{*}\phi^{o})$ as a good generalized parabolic Hitchin pair which in turn gives us the desired torsion-free Hitchin pair on the surface $\mathcal{X}_{T}$.
So now, since we have extended the generic Higgs bundle as a torsion-free Hitchin pair, we obtain the following commutative diagram
[…]
@@
We now choose a suitable connected component $R^{\Lambda,m_{T}}_{S}$ of the atlas of $TFH(\mathcal{X}/S)$, which was constructed in 8.1.1. The component is chosen so that the image of the map $T\longrightarrow TFH(\mathcal{X}/S)$ lies inside the image of the map $R^{\Lambda,m_{T}}_{S}\longrightarrow TFH(\mathcal{X}/S)$. Let $\mathcal{Y}^{H}_{S}$ denote the quot scheme for Gieseker-Hitchin pairs constructed over the quot scheme $R^{\Lambda,m_{T}}_{S}$ (see 8.1.4). Let us denote the projection map by $\tilde{\theta}:\mathcal{Y}^{H}_{S}\longrightarrow R^{\Lambda,m_{T}}_{S}$. Notice that the maps $R^{\Lambda,m_{T}}_{S}\longrightarrow TFH(\mathcal{X}/S)$ and $\mathcal{Y}^{H}_{S}\longrightarrow M^{cl}_{Gie}$ are principal $GL_{N}$-bundles for some large integer $N$ and the map $\tilde{\theta}:\mathcal{Y}^{H}_{S}\longrightarrow R^{\Lambda,m_{T}}_{S}$ is a $GL_{N}$-equivariant morphism (see 8.8).
[…]
@@
Since $T$ is a spectrum of a discrete valuation ring, any $GL_{N}$ bundle on $T$ is trivial. Therefore, by choosing a trivialisation of the $GL_{N}$-bundle $R^{\Lambda,m_{T}}_{S}\times_{TFH(\mathcal{X}/S)}T\to T$ we get the following lift of the diagram 5.8.
[…]
@@
From [1, Proposition 5.11], it follows that the map $\tilde{\theta}:\mathcal{Y}^{H}_{S}\longrightarrow R^{\Lambda,m_{T}}_{S}$ is proper (We discuss the proof below in Lemma 5.4 and Corollary 5.5 for reader’s convenience). So there exists a lift $T\longrightarrow\mathcal{Y}^{H}_{S}$. The composite map $T\longrightarrow\mathcal{Y}^{H}_{S}\longrightarrow M^{cl}_{Gie}$ is our desired extension. ∎
@@
Let $Y\longrightarrow R$ be a quasi-projective morphism of separated schemes over a spectrum of a discrete valuation ring $S$ which is an isomorphism over the generic point of $S$. Suppose any map $T\longrightarrow R$ flat over $S$ can be lifted to a map $T\longrightarrow Y$. Then the map $Y\longrightarrow R$ is proper.
@@
Take the closure of $Y$ inside some projective bundle (over $R$) to get a projective morphism $\overline{Y}\longrightarrow R$. Take any element $y\in\overline{Y}\setminus Y$. Then there exists a map $T\longrightarrow\overline{Y}$ (here, $T$ is a spectrum of a d.v.r) passing through $y$ and flat over $S$. Therefore we get a map $T\longrightarrow R$. But then, by our lifting assumption, this map can be lifted to a map $T\longrightarrow Y$. Let $y^{\prime}\in Y$ be the image of the closed point of $T$ under this map. But, by separateness of the $\overline{Y}$, we must have $y=y^{\prime}$ and therefore $\overline{Y}=Y$, and the map $Y\longrightarrow R$ is proper. ∎
@@
[1, Proposition 5.11] The map $\tilde{\theta}:\mathcal{Y}^{H}_{S}\longrightarrow R^{\Lambda,m_{T}}_{S}$ is proper.
@@
Proof of the claim: Notice that the surface $\mathcal{X}_{T}$ may have a singularity of type $\frac{\mathbb{k}[|x,y,t|]}{(xy-t^{n})}$ because the map $T\to S$ is a faithfully-flat map of spectrum of discrete valuation rings. Let $r_{{}_{T}}:\mathcal{X}^{{res}}_{T}\longrightarrow\mathcal{X}_{T}$ be the minimal resolution of singularities. Then from [18, Proposition 6.5] and [1, §4] it follows that the vector bundle $\mathcal{E}_{T}:=\frac{r_{{}_{T}}^{*}\mathcal{F}_{T}}{Torsion}$ has the property that $(r_{{}_{T}})_{*}\mathcal{E}_{T}\cong\mathcal{F}_{T}$. Notice also that by construction and by the property that $(r_{{}_{T}})_{*}\mathcal{E}_{T}\cong\mathcal{F}_{T}$, it follows that the natural map $r_{{}_{T}}^{*}(r_{{}_{T}})_{*}\mathcal{E}_{T}\longrightarrow\mathcal{E}_{T}$ is a surjective map of sheaves, which means $(\mathcal{E}_{T})|_{R}$ is globally generated. Here $R$ denotes the chain of rational curves in $\mathcal{X}^{{res}}_{T}$. Hence, by definition (Definition 8.1 and Remark 8.2), the vector bundle $\mathcal{E}_{T}$ is a Gieseker vector bundle. The pullback of the Higgs field $\phi_{{}_{T}}$ defines a Higgs field on $\mathcal{E}_{T}$. The pair $(\mathcal{E}_{T},\phi_{T})$ is the desired lift.
@@
6 Flatness of the Hitchin map
In this section, we study the reduced global nilpotent cone of $M^{cl}_{Gie}$ (see 8.1.2), which is the reduced fibre of $h$ over the point $0\in B$. We prove that every irreducible component of of the reduced nilpotent cone has an open subset which is an isotropic substack of $M$ (the derived stack of Higgs bundles) with respect to its log-symplectic form. We use this to compute the dimension of the reduced nilpotent cone and to show that the Hitchin map is flat.
@@ 6 Flatness of the Hitchin map
[…]
The fibre product is usually non-reduced. Let $\mathcal{N}ilp^{red}$ be the reduction of $\mathcal{N}ilp$. It is a stack over the spectrum of a discrete valuation ring $S$. We denote its closed fibre by $\mathcal{N}ilp^{red}_{0}$.
@@
A point in $\mathcal{N}ilp^{red}_{0}$ is a tuple $(X_{r},\mathcal{E},\phi)$, where $X_{r}$ is a Gieseker curve with $r$ $\mathbb{P}^{1}$ bubbles $(0\leq r\leq n)$, $\mathcal{E}$ is a Gieseker vector bundle on $X_{s}$ and $\phi$ is a nilpotent Higgs field, i.e., $\phi^{n}=0$. Given a nilpotent Higgs field, we get a canonical filtration by saturated torsion-free sub-sheaves
$\mathcal{E}^{0}:=0\subsetneq\mathcal{E}^{1}:=\ker\phi\subsetneq\mathcal{E}^{2}:=\ker\phi^{2}\subsetneq\dots\subsetneq\mathcal{E}^{k}:=\ker\phi^{k}\subsetneq\mathcal{E}^{k+1}:=\mathcal{E}$ (6.2)
@@
(Type of a nilpotent torsion-free Higgs pair on the nodal curve $X_{0}$) Let $(\mathcal{F},\psi)$ be a torsion-free Higgs pair on $X_{0}$, where $\mathcal{F}$ is a torsion-free coherent sheaf on $X_{0}$ and $\psi:\mathcal{F}\longrightarrow\mathcal{F}\otimes\omega_{X_{0}}$ is a map of coherent sheaves. Suppose that the Higgs field is nilpotent, i.e., $\psi^{n}=0$, where $n$ denotes the rank of the torsion-free sheaf $\mathcal{F}$. Then, as mentioned above, we get a natural flag of saturated sub sheaves of $\mathcal{F}$.
$\mathcal{F}^{0}:=0\subsetneq\mathcal{F}^{1}:=\ker\psi\subsetneq\mathcal{F}^{2}:=\ker\psi^{2}\subsetneq\dots\subsetneq\mathcal{F}^{k}:=\ker\psi^{k}\subsetneq\mathcal{F}^{k+1}:=\mathcal{F}$ (6.3)
For every $11\leq i\leq k+$ 1, we define $n_{i}:=\textsf{rank}\leavevmode\nobreak\ \left(\frac{\mathcal{F}^{i}}{\mathcal{F}^{i-1}}\right)$ and $d_{i}:=\deg\leavevmode\nobreak\ \left(\frac{\mathcal{F}^{i}}{\mathcal{F}^{i-1}}\right)$, where the degree of a torsion-free sheaf on an irreducible projective curve is defined to be its first Chern number. We say that the nilpotent Higgs bundle $(\mathcal{F},\psi)$ is of type $\{(n_{i},d_{i})\}^{i=k+1}_{i=1}$. We denote it by $\textsf{Type}\leavevmode\nobreak\ (\mathcal{F},\psi)$.
@@
(Type of a nilpotent Higgs bundle on a Gieseker curve $X_{r}$ for some $r\in[0,n]$) Let $r\in[0,n]$ be an integer. Let $\pi_{r}:X_{r}\longrightarrow X_{0}$ be the Gieseker curve with exactly $r$ many $\mathbb{P}^{1}$’s. Let $(\mathcal{E},\phi)$ be a Gieseker-Higgs bundle on $X_{r}$. Then, by definition, $((\pi_{r})_{*}\mathcal{E},(\pi_{r})_{*}\phi)$ is a nilpotent torsion-free Higgs pair on $X_{0}$. We define $\textsf{Type}\leavevmode\nobreak\ (X_{r},\mathcal{E},\phi):=\textsf{Type}\leavevmode\nobreak\ ((\pi_{r})_{*}\mathcal{E},(\pi_{r})_{*}\phi)$.
@@
$\mathcal{E}^{0}:=0\subsetneq\mathcal{E}^{1}:=\ker\phi\subsetneq\mathcal{E}^{2}:=\ker\phi^{2}\subsetneq\dots\subsetneq\mathcal{E}^{k}:=\ker\phi^{k}\subsetneq\mathcal{E}^{k+1}:=\mathcal{E}$ (6.4)
@@
$\mathcal{F}^{0}:=0\subsetneq\mathcal{F}^{1}:=\ker\psi\subsetneq\mathcal{F}^{2}:=\ker\psi^{2}\subsetneq\dots\subsetneq\mathcal{F}^{k}:=\ker\psi^{k}\subsetneq\mathcal{F}^{k+1}:=\mathcal{F}$ (6.5)
Then $(\pi_{r})_{*}\mathcal{E}^{i}\cong\mathcal{F}^{i}$ for every $11\leq i\leq k+$ 1.
@@
Let $(\mathcal{X}^{mod}_{{}_{uni}},\mathcal{E}_{{}_{uni}},\phi_{{}_{uni}})$ denote the restriction of the universal modification, universal vector bundle and the universal Higgs field to $M^{cl}_{Gie}$. We restrict it to $\mathcal{N}ilp^{sm,gen}_{0}$. For simplicity of the notation, let us drop the subscript ”univ” from the notation and simply denote it by $(\mathcal{X}^{mod},\mathcal{E},\phi)$.
@@
$0\longrightarrow\textsf{K}_{c}^{i}\longrightarrow c^{*}\mathcal{E}\xrightarrow{(c^{*}\phi)^{i}}c^{*}\mathcal{E}\otimes\omega^{\otimes i}_{c^{*}\mathcal{X}^{mod}}\longrightarrow\textsf{CK}^{i}_{c}\longrightarrow 0$
@@
$0\longrightarrow\underline{\textsf{K}}^{i}_{c}\longrightarrow(\pi_{r})_{*}(c^{*}\mathcal{E})\xrightarrow{(\pi_{r})_{*}((c^{*}\phi)^{i})=((\pi_{r})_{*}(c^{*}\phi))^{i}}(\pi_{r})_{*}(c^{*}\mathcal{E}\otimes\omega^{\otimes i}_{c^{*}\mathcal{X}^{mod}})\longrightarrow\underline{\textsf{CK}}^{i}_{c}\longrightarrow 0$
Here $\underline{\textsf{K}}^{i}_{c}$ and $\underline{\textsf{CK}}^{i}_{c}$ are kernel and cokernel, respectively. Notice from Lemma 6.5 it follows that $\underline{\textsf{K}}^{i}_{c}=(\pi_{s})_{*}(\textsf{K}^{i}_{c})$. Since the $\textsf{Type}\leavevmode\nobreak\ \leavevmode\nobreak\ (c^{*}{\mathcal{X}^{mod}},c^{*}\mathcal{E},c^{*}\phi)$ is constant over $C$, therefore the Hilbert polynomial of $\underline{\textsf{K}}^{i}_{c}$ does not depend on $c\in C$ which in turn implies that the Hilbert polynomial of $\underline{\textsf{CK}}^{i}_{c}$ does not depend on $c\in C$. Consider next the sheaf cokernel $\underline{\textsf{CK}}^{i}:=Coker((\pi_{r})_{*}((\phi)^{i})$. Since $\underline{\textsf{CK}}^{i}$ is a cokernel, we have that $c^{*}\underline{\textsf{CK}}^{i}=\underline{\textsf{CK}}^{i}_{c}$. Therefore the sheaf $\underline{\textsf{CK}}^{i}$ is flat over $C$. Now consider the kernel sheaf $\underline{\textsf{K}}^{i}:=Ker((\pi_{r})_{*}((\phi)^{i}))$. Since in the four terms sequence
$0\longrightarrow\underline{\textsf{K}}^{i}\longrightarrow(\pi_{r})_{*}(\mathcal{E})\xrightarrow{(\pi_{r})_{*}((\phi)^{i})=((\pi_{r})_{*}(\phi))^{i}}(\pi_{r})_{*}(\mathcal{E}\otimes\omega^{\otimes i}_{\mathcal{X}^{mod}})\longrightarrow\underline{\textsf{CK}}^{i}\longrightarrow 0$
@@
$\mathcal{SC}{(E,\phi)}:=[SC(E,\phi)\xrightarrow{[-,\phi]}SC(E,\phi)\otimes\omega_{X_{r}}]$ (6.6)
where $SC(E,\phi)\subseteq\mathcal{E}ndE$ is the sheaf of local sections $s$ of $\mathcal{E}ndE$ such that $s(E^{i})\subseteq E^{i-1}$ for $i=1,\dots,k+1$.

### v3

@@ 5. Completeness of the Hitchin map
Choose a connected component $R^{\Lambda,m_{T}}_{S}$ of an atlas for $\mathsf{TFH}(\mathcal{X}/S)$ such that the image of $T\to\mathsf{TFH}(\mathcal{X}/S)$ lies in the image of $R^{\Lambda,m_{T}}_{S}\to\mathsf{TFH}(\mathcal{X}/S)$. Let $\mathcal{Y}^{H}_{S}$ be the corresponding Quot scheme of Gieseker-Hitchin pairs over $R^{\Lambda,m_{T}}_{S}$ (see A.3), with projection $\tilde{\theta}\colon\mathcal{Y}^{H}_{S}\to R^{\Lambda,m_{T}}_{S}$. Both maps $R^{\Lambda,m_{T}}_{S}\to\mathsf{TFH}(\mathcal{X}/S)$ and $\mathcal{Y}^{H}_{S}\to M^{cl}_{Gie}$ are principal $GL_{N}$-bundles for sufficiently large $N$, and $\tilde{\theta}$ is $GL_{N}$-equivariant (see A.9).
(5.8) $\mathcal{Y}^{H}_{S}$$M^{cl}_{Gie}$$R^{\Lambda,m_{T}}_{S}$$\mathsf{TFH}(\mathcal{X}/S)$$\scriptstyle{\tilde{\theta}}$$\scriptstyle{GL_{N}-equiv.}$$\scriptstyle{\theta}$
@@
Since $T$ is a spectrum of a discrete valuation ring, any $GL_{N}$ bundle on $T$ is trivial. Therefore, by choosing a trivialisation of the $GL_{N}$-bundle $R^{\Lambda,m_{T}}_{S}\times_{\mathsf{TFH}(\mathcal{X}/S)}T\to T$ we get the following lift of the diagram 5.7.
(5.9) $T^{o}$$\mathcal{Y}^{H}_{S}$$M^{cl}_{Gie}$$T{R^{\Lambda,m_{T}}_{S}}$$\mathsf{TFH}(\mathcal{X}/S)$$\scriptstyle{\tilde{\theta}}$$\scriptstyle{\theta}$
@@
From [2, Proposition 5.11], it follows that the map $\tilde{\theta}:\mathcal{Y}^{H}_{S}\longrightarrow R^{\Lambda,m_{T}}_{S}$ is proper (we discuss the proof below in Lemma 5.5 and Corollary 5.6 for the reader’s convenience). So, there exists a lift $T\longrightarrow\mathcal{Y}^{H}_{S}$. The composite map $T\longrightarrow\mathcal{Y}^{H}_{S}\longrightarrow M^{cl}_{Gie}$ is our desired extension. This completes the proof. ∎
@@
Let $Y\longrightarrow R$ be a quasi-projective morphism of separated schemes over a spectrum of a discrete valuation ring $S$ which is an isomorphism over the generic point of $S$. Assume that both $Y$ and $R$ are flat over $S$. Suppose that any map $T\longrightarrow R$ with $T$ flat over $S$ can be lifted to a map $T\longrightarrow Y$. Then the map $Y\longrightarrow R$ is proper.
@@
Take the closure of $Y$ inside some projective bundle (over $R$) to get a projective morphism $\overline{Y}\longrightarrow R$. Take any element $y\in\overline{Y}\setminus Y$. Then there exists a map $T\longrightarrow\overline{Y}$ (here, $T$ is a spectrum of a d.v.r) passing through $y$ and flat over $S$. Therefore, we get a map $T\longrightarrow R$. But then, by our lifting assumption, this map can be lifted to a map $T\longrightarrow Y$. Let $y^{\prime}\in Y$ be the image of the closed point of $T$ under this map. But, by the separateness of $\overline{Y}$, we must have $y=y^{\prime}$ and therefore $\overline{Y}=Y$, and the map $Y\longrightarrow R$ is proper. ∎
@@
[2, Proposition 5.11] The map $\tilde{\theta}:\mathcal{Y}^{H}_{S}\longrightarrow R^{\Lambda,m_{T}}_{S}$ is proper.
@@
Proof of the claim: Notice that the surface $\mathcal{X}_{T}$ may have a singularity of type $\frac{\mathbb{k}[|x,y,t|]}{(xy-t^{n})}$ because the map $T\to S$ is a faithfully-flat map of spectrum of discrete valuation rings. Let $r_{{}_{T}}:\mathcal{X}^{{res}}_{T}\longrightarrow\mathcal{X}_{T}$ be the minimal resolution of singularities. Then from [22, Proposition 6.5] and [2, §4] it follows that the vector bundle $\mathcal{E}_{T}:=\frac{r_{{}_{T}}^{*}\mathcal{F}_{T}}{Torsion}$ has the property that $(r_{{}_{T}})_{*}\mathcal{E}_{T}\cong\mathcal{F}_{T}$. Notice also that by construction and by the property that $(r_{{}_{T}})_{*}\mathcal{E}_{T}\cong\mathcal{F}_{T}$, it follows that the natural map $r_{{}_{T}}^{*}(r_{{}_{T}})_{*}\mathcal{E}_{T}\longrightarrow\mathcal{E}_{T}$ is a surjective map of sheaves, which means $(\mathcal{E}_{T})|_{R}$ is globally generated. Here $R$ denotes the chain of rational curves in $\mathcal{X}^{{res}}_{T}$. Hence, by definition (Definition A.1 and Remark A.2), the vector bundle $\mathcal{E}_{T}$ is a Gieseker vector bundle. The pullback of the Higgs field $\phi_{{}_{T}}$ defines a Higgs field on $\mathcal{E}_{T}$. The pair $(\mathcal{E}_{T},\phi_{T})$ is the desired lift.
@@
6. Flatness of the Hitchin map
In this section, we study the reduced global nilpotent cone of $M^{cl}_{Gie}$ (see A.1), which is the reduced fibre of $h$ over the point $0\in B$. We prove that every irreducible component of of the reduced nilpotent cone has an open subset which is an isotropic substack of $M$ (the derived stack of Higgs bundles) with respect to its log-symplectic form. We use this to compute the dimension of the reduced nilpotent cone and to show that the Hitchin map is flat.
@@ 6. Flatness of the Hitchin map
(6.1) $\mathcal{N}ilp:=0_{S}\times_{B}M^{cl}_{Gie}$$M^{cl}_{Gie}$$0_{S}$}B{B}
The fibre product is usually non-reduced. Let $\mathcal{N}ilp^{{}^{red}}$ be the reduction of $\mathcal{N}ilp$. It is a stack over the spectrum of a discrete valuation ring $S$. We denote its closed fibre by $\mathcal{N}ilp^{{}^{red}}_{{}_{0}}$.
@@
A point in $\mathcal{N}ilp^{{}^{red}}_{{}_{0}}$ is a tuple $(X_{r},\mathcal{E},\phi)$, where $X_{r}$ is a Gieseker curve with $r$ number of $\mathbb{P}^{1}$ bubbles $(0\leq r\leq n)$, $\mathcal{E}$ is a Gieseker vector bundle on $X_{r}$ and $\phi$ is a nilpotent Higgs field, that is, $\phi^{n}=0$. Given a nilpotent Higgs field, we get a canonical filtration by saturated torsion-free sub-sheaves
(6.2) $\mathcal{E}^{0}:=0\subsetneq\mathcal{E}^{1}:=\ker\phi\subsetneq\mathcal{E}^{2}:=\ker\phi^{2}\subsetneq\dots\subsetneq\mathcal{E}^{k}:=\ker\phi^{k}\subsetneq\mathcal{E}^{k+1}:=\mathcal{E}$
@@
(Type of a nilpotent torsion-free Higgs pair on the nodal curve $X_{0}$) Let $(\mathcal{F},\psi)$ be a torsion-free Higgs pair on $X_{0}$, where $\mathcal{F}$ is a torsion-free coherent sheaf on $X_{0}$ and $\psi:\mathcal{F}\longrightarrow\mathcal{F}\otimes\omega_{X_{0}}$ is a map of coherent sheaves. Suppose that the Higgs field is nilpotent, i.e. $\psi^{n}=0$, where $n$ denotes the rank of the torsion-free sheaf $\mathcal{F}$. Then, as mentioned above, we get a natural flag of saturated sub sheaves of $\mathcal{F}$.
(6.3) $\mathcal{F}^{0}:=0\subsetneq\mathcal{F}^{1}:=\ker\psi\subsetneq\mathcal{F}^{2}:=\ker\psi^{2}\subsetneq\dots\subsetneq\mathcal{F}^{k}:=\ker\psi^{k}\subsetneq\mathcal{F}^{k+1}:=\mathcal{F}$
For every $1\leq i\leq k+1$, we define $n_{i}:=\textsf{rank}\penalty 10000\ \left(\frac{\mathcal{F}^{i}}{\mathcal{F}^{i-1}}\right)$ and $d_{i}:=\deg\penalty 10000\ \left(\frac{\mathcal{F}^{i}}{\mathcal{F}^{i-1}}\right)$, where the degree of a torsion-free sheaf on an irreducible projective curve is defined to be its first Chern number. We say that the nilpotent Higgs bundle $(\mathcal{F},\psi)$ is of type $\{(n_{i},d_{i})\}^{i=k+1}_{i=1}$. We denote it by $\textsf{Type}\penalty 10000\ (\mathcal{F},\psi)$.
@@
(Type of a nilpotent Higgs bundle on a Gieseker curve $X_{r}$ for some $r\in[0,n]$) Let $r\in[0,n]$ be an integer. Let $\pi_{r}:X_{r}\longrightarrow X_{0}$ be the Gieseker curve with exactly $r$ many $\mathbb{P}^{1}$’s. Let $(\mathcal{E},\phi)$ be a Gieseker-Higgs bundle on $X_{r}$. Then, by definition, $((\pi_{r})_{*}\mathcal{E},(\pi_{r})_{*}\phi)$ is a nilpotent torsion-free Higgs pair on $X_{0}$. We define $\textsf{Type}\penalty 10000\ (X_{r},\mathcal{E},\phi):=\textsf{Type}\penalty 10000\ ((\pi_{r})_{*}\mathcal{E},(\pi_{r})_{*}\phi)$.
@@
(6.4) $\mathcal{E}^{0}:=0\subsetneq\mathcal{E}^{1}:=\ker\phi\subsetneq\mathcal{E}^{2}:=\ker\phi^{2}\subsetneq\dots\subsetneq\mathcal{E}^{k}:=\ker\phi^{k}\subsetneq\mathcal{E}^{k+1}:=\mathcal{E}$
@@
(6.5) $\mathcal{F}^{0}:=0\subsetneq\mathcal{F}^{1}:=\ker\psi\subsetneq\mathcal{F}^{2}:=\ker\psi^{2}\subsetneq\dots\subsetneq\mathcal{F}^{k}:=\ker\psi^{k}\subsetneq\mathcal{F}^{k+1}:=\mathcal{F}$
Then $(\pi_{r})_{*}\mathcal{E}^{i}\cong\mathcal{F}^{i}$ for every $1\leq i\leq k+1$.
@@
Let $(\mathfrak{X}_{{}_{uni}},\mathcal{E}_{{}_{uni}},\phi_{{}_{uni}})$ denote the restriction of the universal modification, universal vector bundle and the universal Higgs field to $M^{cl}_{Gie}$. We restrict it to $\mathcal{N}ilp^{sm,gen}_{0}$. For simplicity of the notation, let us drop the subscript ”univ” from the notation and simply denote it by $(\mathfrak{X},\mathcal{E},\phi)$.
@@
$0\longrightarrow\textsf{K}_{c}^{i}\longrightarrow c^{*}\mathcal{E}\xrightarrow{(c^{*}\phi)^{i}}c^{*}\mathcal{E}\otimes\omega^{\otimes i}_{c^{*}\mathfrak{X}}\longrightarrow\textsf{CK}^{i}_{c}\longrightarrow 0$
@@
$0\longrightarrow\underline{\textsf{K}}^{i}_{c}\longrightarrow(\pi_{r})_{*}(c^{*}\mathcal{E})\xrightarrow{(\pi_{r})_{*}((c^{*}\phi)^{i})=((\pi_{r})_{*}(c^{*}\phi))^{i}}(\pi_{r})_{*}(c^{*}\mathcal{E}\otimes\omega^{\otimes i}_{c^{*}\mathfrak{X}})\longrightarrow\underline{\textsf{CK}}^{i}_{c}\longrightarrow 0$
Here $\underline{\textsf{K}}^{i}_{c}$ and $\underline{\textsf{CK}}^{i}_{c}$ are kernel and cokernel, respectively. Notice from Lemma 6.5 it follows that $\underline{\textsf{K}}^{i}_{c}=(\pi_{s})_{*}(\textsf{K}^{i}_{c})$. Since the $\textsf{Type}\penalty 10000\ \penalty 10000\ (c^{*}{\mathfrak{X}},c^{*}\mathcal{E},c^{*}\phi)$ is constant over $C$, therefore the Hilbert polynomial of $\underline{\textsf{K}}^{i}_{c}$ does not depend on $c\in C$ which in turn implies that the Hilbert polynomial of $\underline{\textsf{CK}}^{i}_{c}$ does not depend on $c\in C$. Consider next the sheaf cokernel $\underline{\textsf{CK}}^{i}:=Coker((\pi_{r})_{*}((\phi)^{i})$. Since $\underline{\textsf{CK}}^{i}$ is a cokernel, we have that $c^{*}\underline{\textsf{CK}}^{i}=\underline{\textsf{CK}}^{i}_{c}$. Therefore, the sheaf $\underline{\textsf{CK}}^{i}$ is flat over $C$. Now consider the kernel sheaf $\underline{\textsf{K}}^{i}:=Ker((\pi_{r})_{*}((\phi)^{i}))$. Since in the four terms sequence
$0\longrightarrow\underline{\textsf{K}}^{i}\longrightarrow(\pi_{r})_{*}(\mathcal{E})\xrightarrow{(\pi_{r})_{*}((\phi)^{i})=((\pi_{r})_{*}(\phi))^{i}}(\pi_{r})_{*}(\mathcal{E}\otimes\omega^{\otimes i}_{\mathfrak{X}})\longrightarrow\underline{\textsf{CK}}^{i}\longrightarrow 0$
@@
(6.6) $\mathcal{SC}{(E,\phi)}:=[SC(E,\phi)\xrightarrow{[-,\phi]}SC(E,\phi)\otimes\omega_{X_{r}}]$
where $SC(E,\phi)\subseteq\mathcal{E}\mathsf{nd}\penalty 10000\ E$ is the sheaf of local sections $s$ of $\mathcal{E}\mathsf{nd}\penalty 10000\ E$ such that $s(E^{i})\subseteq E^{i-1}$ for $i=1,\dots,k+1$.